OpenStudy (andromeda992):

Please help! Is my work correct? If the velocity of an object changes from 65 m/s to 98m/s during a time interval of 12 s, what is the acceleration of the object? A. 33 m/s B. 5.42 m/s C. 2.75 m/s D. 13.58 m/s

OpenStudy (andromeda992):

I think that C is the answer

OpenStudy (andromeda992):

\[a=\frac{ 98-65 }{ 12 } = \frac{ 33 }{ 12 } =2.75\]

OpenStudy (andromeda992):

Is that correct?

OpenStudy (kgrendel0324):

Yes you are :D

OpenStudy (moldybubblegum12):

@Andromeda992 You're correct

OpenStudy (andromeda992):

Thank y'all

OpenStudy (andromeda992):

*thanks

OpenStudy (mathstudent55):

You are correct. Acceleration is the rate of change of speed. \(a = \dfrac{change~in~speed}{time} =\) \(\Large = \dfrac{98 \frac{m}{s} -65 \frac{m}{s} }{12~s} \) \(\Large = 2.75 \dfrac{m}{s^2} \) Notice the unit of acceleration is \(\dfrac{m}{s^2} \), not \(\dfrac{m}{s} \).

OpenStudy (andromeda992):

Thanks

OpenStudy (mathstudent55):

You're welcome.